### v1

# SurgVidLM: Towards Multi-grained Surgical Video Understanding with Large Language Model

Abstract
Recent advances in Multimodal Large Language Models have demonstrated great potential in the medical domain, facilitating users to understand surgical scenes and procedures. Beyond image-based methods, the exploration of Video Large Language Models (Vid-LLMs) has emerged as a promising avenue for capturing the complex sequences of information involved in surgery. However, there is still a lack of Vid-LLMs specialized for fine-grained surgical video understanding tasks, which is crucial for analyzing specific processes or details within a surgical procedure. To bridge this gap, we propose SurgVidLM, the first video language model designed to address both full and fine-grained surgical video comprehension. To train our SurgVidLM, we construct the SVU-31K dataset which consists of over 31K video-instruction pairs, enabling both holistic understanding and detailed analysis of surgical procedures. Furthermore, we introduce the StageFocus mechanism which is a two-stage framework performing the multi-grained, progressive understanding of surgical videos. We also develop the Multi-frequency Fusion Attention to effectively integrate low and high-frequency visual tokens, ensuring the retention of critical information. Experimental results demonstrate that SurgVidLM significantly outperforms state-of-the-art Vid-LLMs in both full and fine-grained video understanding tasks, showcasing its superior capability in capturing complex procedural contexts.
Keywords:
Surgical Video Understanding Video Language Model Multi-grained Analysis.
1 Introduction
Surgical visual-language models have attracted considerable attention due to their advancements in surgical scene understanding, encompassing tasks such as anatomy recognition [31, 21], instrument detection [2, 3], and workflow recognition [25, 17]. These techniques demonstrate substantial potential in enhancing surgical training and medical diagnosis. Meanwhile, the development of Multimodal Large Language Models (MLLMs) has led to various studies utilizing multimodal data from surgical procedures to perform image-level surgical scene understanding [19, 8, 9, 23, 16, 20]. Trained through medical datasets, MLLMs can query biomedical images and provide more flexible responses to medical inquiries based on their inherent reasoning ability. However, image-based surgical scene understanding methods primarily focus on static visual cues, limiting their ability to capture the complex temporal dependencies and procedural context inherent in surgical videos. To address this limitation, Video Large Language Models (Vid-LLMs) extend the reasoning and interaction capabilities of LLMs to video understanding by using parallel video-text datasets [26]. By extracting temporal features, Vid-LLM can effectively perform joint modeling of temporal representations and text [1, 6]. However, the acquisition of video-text datasets remains a significant challenge, particularly in the surgical domain, due to the requirement for costly expert annotations and specialized knowledge. Despite efforts to construct of surgical video understanding data, existing datasets [10, 8] are limited by accessibility or lack of reasoning annotations.
Recently, extensive progress has been made on video understanding in computer vision field [4, 30, 12, 28, 11]. However, these model face challenges when directly applied to surgical domain due to their reliance on general video datasets, which differ significantly from the complex, domain-specific context inherent in surgical procedures [5]. In the surgical scenario, LLaVA-Surg [10] proposes the Surg-QA dataset, comprising 102K video-instruction pairs. However, Surg-QA is not publicly available. Although Surgical-LLaVA [8] achieves the integration of image and video visual representations for video understanding, it is fine-tuned using image-based annotations. Moreover, existing studies only focus on full video understanding without delving into fine-grained video tasks. In this case, models can hardly analyze a specific process or detail within a surgical procedure. Due to computational resource constraints, Vid-LLMs extract visual information by low-frequency sampling of video frames, which inevitably results in the loss of critical local information essential for fine-grained video understanding [15, 29]. If Vid-LLMs can locate the time of video clips related to questions and sample frames at a higher frequency specifically, the visual information will be better extracted, thereby facilitating more precise fine-grained analysis.
To address these challenges, we first construct SVU-31K, a large-scale surgical video dataset containing both full video and fine-grained video understanding annotations. Based on SVU-31K, we propose the SurgVidLM, the first surgical video language model addressing the gap in multi-grained surgical video comprehension. Specifically, our contributions are as follows: (1) We create SVU-31K, a large-scale multi-grained surgical video understanding dataset annotated through a novel knowledge augmentation pipeline. SVU-31K designs multiple fine-grained surgical video understanding tasks specific to procedure interaction and perception. It also contains annotations of full videos, which allow the model to provide context information for fine-grained understanding. (2) We propose SurgVidLM, a two-stage Vid-LLM adopting the StageFocus mechanism for progressive multi-grained surgical video understanding. Additionally, Multi-frequency Fusion Attention is integrated to facilitate the interaction of multi-grained visual tokens. (3) Extensive comparison and ablation studies prove the outstanding performance of our SurgVidLM, demonstrating its potential in multi-grained surgical video understanding.
Figure 1: Data collection and construction pipeline for SVU-31K.
2 Methodology
2.1 Proposed Dataset: Surgical Video Understanding (SVU)-31K
Beyond the scarcity of benchmarks for surgical video understanding, existing studies [10, 18] do not investigate fine-grained video understanding. To bridge this gap, we present SVU-31K, a high-quality instruction-tuning dataset tailored to endow Vid-LLMs to understand multi-grained surgical videos. The construction pipeline comprises the following three steps, which are shown in Fig. 1.
2.1.1 Video Search and Preprocess.
We collect 1882 surgical procedural videos from YouTube and WebSurg, encompassing a range of surgical specialties: Pancreatic Surgery, Digestive Surgery, Hernia Surgery, and Colorectal, etc. Each video is accompanied by its abstract text, as well as keystep annotations with corresponding timecodes. To structure the data, we segment each video into clips based on the annotated key steps. A pre-trained ResNet-50 model is then employed to filter out non-surgical segments, ensuring that only clips containing surgical content are retained. Specifically, we annotate surgical images in 200 videos and split them into an 80/20 train-test set. The trained model achieves 94% classification accuracy. Then, we sample at 1 fps from each segment and classify the segment as surgical if at least 80% of the frames are predicted as surgical—this yields 100% accuracy for the segment filter. The audio from the remaining clips is transcribed into narrations using OpenAI Whisper [27], using LLM for precise transcription of medical terminology. As a result, we extract a three-level hierarchical knowledge representation comprising Abstract, Keystep, and Narration, which serves as a structured foundation for the following instruction-tuning.
2.1.2 Multi-grained Knowledge Augmentation Through GPT-4.
While texts at varying levels provide diverse insights into surgical videos, there are still some limitations. Specifically, Abstracts often contain redundant content. Keysteps, while concise and structured, are limited in linguistic diversity and lack detailed contextual information. Meanwhile, narratives, extracted from audio transcriptions, tend to be fragmented and incoherent. These incomplete, noisy, or overly concise textual inputs can lead to suboptimal question generation, ultimately limiting the model’s ability to reason about surgical procedures effectively. To address this, we introduce multi-grained knowledge augmentation to enhance textual quality across all levels. First, we employ GPT-4 to reorganize and supplement the abstract texts based on keysteps, removing extraneous details. Then, keystep texts are enriched with narrative details, adding procedural, anatomical, and instrumental context. We integrate narration texts into the keystep texts rather than expanding on them, as such clips are too brief to be used directly.
2.1.3 Multi-grained Instruction-Tuning Data.
Using the augmented texts as context, we construct a multi-grained instruction-tuning dataset with GPT-4, introducing granularity-specific visual question-answer pairs to enhance the model’s understanding of surgical videos. At the full-video level, the model is trained on summarization and perception tasks. For keystep-based video clips, we design three types of video understanding tasks, as illustrated in the purple part of Fig. 1: (i) Summary description task requires the model to generate a concise summary of surgical keysteps. (ii) Temporal visual reasoning task (Plan, Pre-requisites) intends to understand the temporal order of surgical operations and infer the prerequisites or planning steps for specific actions. For example, if a clip shows the suturing process, the task might require the model to deduce that the prerequisite steps of tissue alignment and needle preparation are needed prior to suturing. (iii) Perceptual visual reasoning task (Causal, Notice and Suggestion, Recall) infers about causal relationships, suggestions, and recalls key information during surgical procedures. When observing a surgeon encountering difficulty with tissue manipulation, the model should analyze the causal relationship between improper force application and tissue damage, suggest a correction, and recall prior successful approaches in similar scenarios. Each keystep-based QA pair is associated with a timecode, enabling the model to align the QA pair with a specific video clip when extracting visual information. Finally, the SVU-31K dataset includes 1.8K QA pairs for the full video and 29.4K QA pairs for fine-grained video understanding tasks.
2.2 Proposed Model: SurgVidLM
2.2.1 Preliminaries
We develop the SurgVidLM upon the Qwen2-VL framework, which integrates vision encoders and language model Qwen2[24]. Let $V=\{V(t)\in\mathbb{R}^{H\times W\times 3}\mid t=0,\dots,T\}$ represents a video sequence comprising $T$ frames, along with an associated query $q$. The visual encoder extracts frame-level features from downsampled frames $V_{F}$ which are sparsely sampled at fixed intervals $I_{F}$, yielding low-frequency visual tokens $X_{F}$. Subsequently, $X_{F}$ is aligned with language tokens and processed by the LLM to generate the response.
Figure 2: The architecture of SurgVidLM. Stage 1 focuses on a holistic understanding of the full video, while Stage 2 fuses the information from both the current and previous stages to achieve accurate, fine-grained video understanding.
2.2.2 StageFocus for Multi-grained Video Understanding.
[…]
2.2.3 Multi-frequency Fusion Attention.
Due to the inherent coarseness of the low-frequency visual tokens and the detailed specificity of high-frequency visual tokens, directly concatenating them for input into the language model may result in suboptimal performance [15]. Effective fine-grained video understanding requires capturing the interactions among multiple events, necessitating the integration of global context with local features. Therefore, we design the Multi-frequency Fusion Attention (MFA):
$E\left(X_{C},X_{F}\right)=softmax\left(\frac{X_{C}X_{F}^{T}}{\sqrt{d}}\right)X_{C}$ (1)
where we normalize the attention weights, ensuring that each high-frequency token is informed by the global context of the low-frequency tokens. This integration enhances the model’s ability to retain the relevance of local details while preserving a broader global context.
2.2.4 Training Strategy.
Following the StageFocus mechanism, we design a two-phase training strategy that endowed SurgVidLM with the capacity for multi-grained surgical video understanding. The first phase is dedicated to knowledge augmentation, where the model is trained for a holistic understanding of long surgical video sequences. In this phase, we utilize QA pairs corresponding to full videos from the SVU-31K dataset. The language model (LLM) is fine-tuned using LoRA [7]. The second phase focuses on enhancing fine-grained surgical video understanding and MFA integration. Under the StageFocus mechanism, we train the LLM and MFA modules with the fine-grained task part of the SVU-31K dataset, facilitating cross-stage information fusion. The visual encoder’s parameters remain frozen in both stages.
3 Experiments
3.1 Implementation Details
We adopt the pretrained Qwen2-VL-72B-Instruct [22] as the initial weights of our SurgVidLM. In the first training phase, we use a batch size of 4, while in the second phase, the batch size is increased to 16. The learning rate is set to 1e-5, and the training spans a total of 3 epochs. The resolution of videos is 224 $\times$ 224. The training process is conducted on 6 NVIDIA H100 GPUs, with the first phase taking approximately 1.5 hours and the second phase requiring around 32 hours. For comparison, we evaluate SurgVidLM against Video-LLaVA [13], GroundingGPT [12], Video-LLaMA-3 [28], and Qwen2-VL [22]. To provide a thorough evaluation of video description, we follow the evaluation pipeline in  [14], employing the GPT-4 model to assess the responses on a scale from 1 to 5 across four key aspects: correctness of information, detail orientation, contextual understanding, and Temporal Understanding. For the reasoning tasks, we utilize standard quantitative metrics, including BLEU-4, METEOR, ROUGE-L (scaled from 0 to 1 and presented as percentages in tables), and CIDEr (scaled from 1 to 10).
Table 1: Comparison of SurgVidLM with zero-shot Vid-LLMs across multi-grained video understanding tasks on the SVU-31K dataset. CI, DO, CU and TU means Correctness of Information, Detail Orientation, Contextual Understanding and Temporal Understanding, respectively.
Model Full Video Description Fine-grained Video Description
CI DO CU TU CI DO CU TU
Video-LLaVA 1.03 1.01 1.01 1.03 1.15 1.09 1.24 1.21
GroundingGPT 1.05 1.06 1.15 1.08 1.05 1.16 1.26 1.16
Video-LLaMA-3 1.43 1.29 1.66 1.63 1.65 1.62 2.16 1.95
Qwen2-VL 1.22 1.09 1.64 1.53 1.82 1.77 2.49 1.95
Ours 2.83 2.52 3.09 2.62 2.83 3.08 3.51 2.86
Model Fine-grained Temporal Visual Reasoning Fine-grained Perception Visual Reasoning
BLEU-4 CIDEr METEOR ROUGE-L BLEU-4 CIDEr METEOR ROUGE-L
Video-LLaVA 5.97 2.29 14.02 26.19 10.01 5.51 27.88 29.02
GroundingGPT 2.80 1.44 13.52 30.86 4.47 0.51 39.55 35.34
Video-LLaMA-3 2.94 0.82 12.45 12.66 6.18 1.89 16.95 16.99
Qwen2-VL 5.67 1.79 10.06 25.96 8.51 2.39 20.51 18.93
Ours 11.33 3.59 16.37 32.74 16.19 9.01 41.88 37.75
3.2 Results
Table 1 presents the qualitative comparison between SurgVidLM and several SOTA Vid-LLMs. For full-video descriptions, our model significantly outperforms SOTA approaches across all four metrics, with improvements of at least 1.4, 1.23, 1.34, and 0.99, demonstrating the effectiveness of the instruction-tuning process on the SVU-31K dataset. Similarly, for the fine-grained video description, our model also leads the field, outperforming all baselines by 1.01, 1.31, 1.02, and 0.91, highlighting its precision in describing complex surgical scenes.
We evaluate spatiotemporal reasoning in surgical videos, focusing on temporal and perceptual visual reasoning. For temporal reasoning, SurgVidLM outperforms all baselines, achieving gains of 5.36, 0.13, 2.35, and 1.88 across BLEU-4, CIDEr, METEOR, and ROUGE-L. This success is driven by Multi-grained Fusion Attention, enabling video analysis at multiple granularities and effective temporal reasoning. It also leverages our dataset’s Plan and Pre-requisites tasks to infer procedural order (e.g., tissue alignment before suturing). For perceptual reasoning, SurgVidLM surpasses Qwen2-VL, benefiting from the StageFocus mechanism and instruction-tuning tasks like Causal and Recall.
Fig. 3 highlights SurgVidLM’s superior procedural reasoning and anatomical understanding, accurately identifying key structures (e.g., Denonvilliers’ fascia) and surgical rationales (e.g., incision placement above the Douglas pouch). Unlike Video-LLaMA-3 and Qwen2-VL, which give generic responses, SurgVidLM provides step-wise surgical objectives (mesorectal excision) and risks (avoiding tissue injury), demonstrating the effectiveness of instruction-tuning tasks like Plan, Pre-requisites, Causal, Notice, and Suggestion.
Figure 3: Qualitative examples of fine-grained video reasoning tasks.
Table 2: Ablation study on StageFoucs and Multi-grained Fusion Attention.
StageFocus Multi-grained Fusion Attention Fine-grained Video Description Fine-grained Visual Reasoning
CI DO CU TU BLEU-4 CIDEr METEOR ROUGE-L
2.43 2.85 3.22 2.57 0.108 0.590 0.351 0.359
2.57 2.95 3.39 2.77 0.126 0.782 0.371 0.384
2.83 3.08 3.51 2.86 0.161 0.901 0.418 0.377
Our ablation study confirms the complementary roles of StageFocus and Multi-frequency Fusion Attention, with their combination achieving the best performance across video description and visual reasoning tasks (Table 2). Using StageFocus alone reduces the correctness of information score by 0.26 in fine-grained video description, highlighting the importance of MFA for integrating global visual context. Meanwhile, StageFocus effectively preserves detailed and contextual information. These results demonstrate that both mechanisms work together to enhance accuracy and completeness.
4 Conclusions
In this work, we introduce SVU-31K, a large-scale multi-grained surgical video understanding dataset that comprises over 31K video-instruction pairs. Based on SVU-31K, we propose SurgVidLM, the first surgical video language model specifically designed for multi-grained surgical video comprehension. Our proposed StageFocus mechanism utilizes a two-stage reasoning approach to guide the model to analyze surgical procedures holistic to detail progressively. The integration of Multi-frequency Fusion Attention further enhances the model’s performance by effectively combining low- and high-frequency visual tokens, thus preserving critical information essential for fine-grained surgical scene understanding. Experimental results demonstrate that SurgVidLM significantly outperforms SOTA Vid-LLMs across various multi-grained surgical video understanding tasks. These findings underscore the potential of SurgVidLM to advance the field of surgical video understanding. Future work will focus on expanding the dataset and exploring the application of SurgVidLM in real-time surgical environments, thereby contributing to improved clinical outcomes through enhanced video analysis capabilities.